# Poker Hand History File Format Specification

### 18 Dec 2023 v1

@@ Abstract
Board games, Card games, Games of chance, Game design, Multi-agent systems, Poker, Rule based systems, Scripting, Strategy games
@@ I Introduction
Pioneering breakthroughs in imperfect information game-solving algorithms have often been accompanied by their applications in state-of-the-art poker AI agents as seen in the example of Rhode Island hold ’em [1], fixed-limit Texas hold ’em [2], heads-up no-limit Texas hold ’em [3, 4], and 6-max no-limit Texas hold ’em [5]. Successive state-of-the-art algorithms were shown to solve games of progressively larger game tree sizes.
@@ II Motivational Examples
  "Bryce Yockey",
  "Phil Hui",
  "John Esposito",
  "Josh Arieh",
@@ III Related Works
Table ’?????’ ?-max Seat #3 is the button
@@
Figure 4: A sample hand played by Pluribus in the supplementary of Brown and Sandholm [5]. Note that the line was broken in multiple places to fit in a single column.
@@ V-A2 Antes
V-A3 Blinds or Straddles
The “blinds_or_straddles” field, representing all forced bets except the bring-in, must only be specified in button games as an array of non-negative integers or floats of length equal to the number of players. Most button games have two blinds: a small blind and a big blind. These are simply described by an array with two positive values in the first two indices while all other elements are zero. Optionally, various straddle regimes ranging from under-the-gun straddle, double straddle, button straddle, et cetera, can be introduced by setting the elements in the relevant indices with non-zero values. Samples are demonstrated in Figure 6. Note that, in heads-up situations, the effective blind amounts are the reversed version of the array field value.
# Small and big blind
@@ V-A3 Blinds or Straddles
V-A4 Bring-in
The “bring_in” field is of a positive integer or float value and is primarily used in stud poker variants to represent the bring-in forced bet. The usage of bring-ins is mutually exclusive with blinds or straddles. In other words, both must never be specified together. A sample definition is shown in Figure 7.
@@ V-B1 Actions
  • Dealing board cards and
@@ V-C Action Notation
Note that commentaries and TOML comments are two distinct mechanisms in the PHH format. They are both followed by a single hash symbol. However, TOML comments are ignored by the parser, while the commentary information is considered a part of the action notation by the parsing system. The annotator should use commentaries to give extra contextual information about the hand that cannot be determined by the parser such as a summary of the metagame between some of the players, exposed hole cards, and interruptions, just to name a few, whereas the TOML comments should serve as a guide to the reader who is directly reading the file without the help of a software parser. The information provided by TOML comments could be the names of the betting round or the actor.
@@ V-C5 Folding
is the folding action. The action samples are shown in Figures 1 and 2.
V-C6 Checking or Calling
@@
V-C7 Completion, Betting, or Raising to
@@
V-C8 Showing or Mucking Hole Cards
@@ VI-E3 Currency
VI-E4 Ante Trimming Status
@@
VI-F Time Control
VI-F1 Time Limit
@@ VI-F2 Time Banks
VII User-defined Fields

### 01 Jan 2024 v2

@@ Abstract
Card games, Games of chance, Game design, Multi-agent systems, Poker, Rule based systems, Scripting, Strategy games
@@ I Introduction
Pioneering breakthroughs in imperfect information game-solving algorithms have often been accompanied by their applications in state-of-the-art poker AI agents as seen in the example of Rhode Island hold ’em [1], fixed-limit Texas hold ’em [2], heads-up no-limit Texas hold ’em [3, 4], and 6-player no-limit Texas hold ’em [5]. Successive state-of-the-art algorithms were shown to solve games of progressively larger game tree sizes.
@@ II Motivational Examples
  "Bryce Yockey", "Phil Hui",
  "John Esposito", "Josh Arieh",
@@ III Related Works
Table ‘?????’ ?-max Seat #3 is the button
@@
Figure 4: A sample hand played by Pluribus in the supplementary of Brown and Sandholm [5]. Note that the line was broken to fit in the page.
@@ VI-A2 Antes
VI-A3 Blinds or Straddles
The “blinds_or_straddles” field, representing all forced bets except the bring-in, must only be specified in button games as an array of non-negative integers or floats of length equal to the number of players. Most button games have two blinds: a small blind and a big blind. These are simply described by an array with two positive values in the first two indices while all other elements are zero. Optionally, various straddle regimes ranging from under-the-gun straddle, double straddle, button straddle, et cetera, can be introduced by setting the elements in the relevant indices with non-zero values. Examples are demonstrated in Figure 6. Note that, in heads-up situations, the effective blind amounts are the reversed version of the array field value.
# Small and big blind (heads-up)
@@ VI-A3 Blinds or Straddles
VI-A4 Bring-in
The “bring_in” field is of a positive integer or float value and is primarily used in stud poker variants to represent the bring-in forced bet. The usage of bring-ins is mutually exclusive with blinds or straddles. In other words, both must never be defined together. A sample definition is shown in Figure 7.
@@ VI-B1 Actions
  • Dealing board cards or
@@ VI-C Action Notation
Note that commentaries and TOML comments are two distinct mechanisms in the PHH format. They are both followed by a single hash symbol. However, TOML comments are ignored by the parser, while the commentary information is considered a part of the action notation by the parsing system. The annotator should use commentaries to give extra contextual information about the hand that cannot be determined by the parser such as a summary of the metagame between some of the players, exposed hole cards, and interruptions, just to name a few, whereas the TOML comments should serve as a guide to the reader who is directly reading the file without the help of a software parser. For instance, the information provided by TOML comments may include the names of the betting round or the actor.
@@ VI-C5 Folding
is the folding action. Examples are shown in Figures 1 and 2.
VI-C6 Checking or Calling
@@
VI-C7 Completion, Betting, or Raising to
@@
VI-C8 Showing or Mucking Hole Cards
@@ VII-E3 Currency
VII-E4 Ante Trimming Status
@@
VII-F Time Control
VII-F1 Time Limit
@@ VII-F2 Time Banks
VIII User-defined Fields